Part 1. The tide gate opens automatically when the tidewater at B subsides, allowing the marsh at A to drain. For the water level h = 4 m, determine the horizontal reaction at the smooth stop C. The gate has a width of 2 m. At what height h will the gate be on the verge of opening?


Part 2. Determine the critical height h of the water level where the concrete gravity darn starts to tip over. The density of concrete is pc = 240 Mg/m3. Hint: Work the problem using a 1-m width of the dam.



1
Expert's answer
2021-11-21T17:27:43-0500

Since the gate has a constant width of b = 2 m, the intensities of the distributed load on the left and right sides of the gate at C are

"(W_{C})_{L} = \rho _{w}gh_{BC} (b)
= 78.48\cdot 10^{3}~N/m,"

"(W _\nC\n\u200b\n ) \n_R\n\u200b\n =\u03c1 \n_w\n\u200b\n gh _{\nAC}\n\u200b\n (b)\n= 68.67\\cdot 10^{3}~ N\/m,"

the resultant triangular distributed load on the left and right sides of the gate is shown on its free-body diagram, Fig. a,

"FL = \frac{1}{2}(W_{C})_{L}L_{BC} = 156.96\cdot10^{3}~ N,








\\
FR = \frac{1}{2} (W_{C})_{L}L_{AC} =120.17\cdot 10^{3}~ N,"

referring to the free-body diagram of the gate in Fig. a,

"\Sigma M_D=0,"

"F_C=25.3~kN."